1- The concentration of bromide ion may be determined by gravimetric analysis, using this reaction:
Ag * (aq) Br (aq) AgBr (s)
A 0.500-L aqueous solution containing bromide was reacted with excess Ag. If the AgBr precipitate formed in this reaction has a mass of 0.0035 g, how many moles of Br were in the solution? What was the molarity of Br in the solution?

2-Identify each of these reactions as precipitation, neutraliza tion, or metal displacement:
a. Ba (OH), (aq) + FeCl2 (5) - BaCl (aq) + Fe (OH)(s)
b. Ba (OH), (aq) + 2HCl (aq) - BaCl (aq) + 2H2O

c. 3 Ba (s) + 2 Al (NO3) 3 (aq) 3 Ba (NO3), (aq) + 2 Al (s)
Answer #1

Molar mass of Br = 79.91 g/mol

Molar mass of AgBr =107.87 + 79.91 = 187.78 g/mol

Consider precipitation reaction, Ag + (aq) + Br - (aq) \rightarrow AgBr (s)

From above reaction, We get relation 1 mol AgBr \equiv 1 moles Br -

187.78 g AgBr \equiv 79.91 g Br -

\therefore 0.0035 g AgBr \equiv 79.91 \times 0.0035 / 187.78 g Br -

0.0035 g AgBr \equiv 0.00149 g Br -

\therefore No. of moles of Br - = Mass of Br - / Molar mass of Br - = 0.00149 g / 79.91 g/mol = 1.86 \times 10 -05 mol  

1.86 \times 10 -05 mol  Br - is present in 0.500 L aqueous solution.

We have, Molarity = No.of moles of solute / Volume of solution in L

[Br -] = (1.86 \times 10 -05 mol) / 0.500 L = 3.7  \times 10 -05 mol

ANSWER : [Br -] = 3.7  \times 10 -05 mol

PART 2

Precipitation reaction : It is the reaction in which one of the product is solid.

Neutralization reaction : It is the reaction between acid and base to produce salt and water.

Metal displacement reaction: In this reaction, metal ion from reactant compound is displaced by another ion.

Reaction observation Conclusion
Ba(OH) 2 (aq) + FeCl 2 (aq) \rightarrow BaCl 2 (aq) + Fe(OH) 3 (s) Product Fe(OH) 3 is solid. Precipitation reaction.
Ba(OH) 2 (aq) + 2 HCl (aq) \rightarrow  BaCl 2 (aq) + 2 H2O (l) base and acid reacts and produce salt and water Acid-Base reaction
3 Ba (S) + 2 Al (NO3) 3 (aq) \rightarrow2  Ba(NO3) 2 + Al 3+ (Aq) Metal ion Al 3+ is displaced from reactant aluminium nitrate by Ba 2+ Metal displacement reaction